# Forum for Information Retrieval Evaluation, December 12-15, 2024, India

### v1

@@ 1 Introduction
Sarcasm, a form of expression where the intended meaning sharply diverges from the literal meaning, poses a significant challenge for sentiment analysis systems [1, 2, 3]. This complexity arises because sarcasm often involves subtle linguistic cues and context-dependent interpretation, making it difficult for automated systems to accurately detect and classify [4, 5, 6, 7, 8].
@@
Code-mixing [9, 10, 11] in NLP refers to the blending of two or more languages within a single sentence or conversation, commonly seen in multilingual societies. It poses unique challenges for tasks like language identification, part-of-speech tagging, and sentiment analysis. Traditional NLP models often struggle with code-mixed data due to the complex interaction between languages. To address this, researchers have developed advanced neural network models, such as RNNs and CNNs, that better capture linguistic context. Code-mixed text also complicates machine translation and speech recognition, requiring specialized approaches to handle language switching effectively. As multilingualism becomes more prevalent, the importance of developing robust NLP systems capable of processing code-mixed data continues to grow.
Code-mixing is a widespread phenomenon in multilingual societies, particularly in Dravidian languages such as Tamil and Malayalam, where users frequently switch between native languages and English [12, 9, 13]. These code-mixed texts often appear in non-native scripts, such as Roman, due to the convenience of typing on digital platforms [9, 14, 15]. This creates a unique set of challenges for natural language processing (NLP) systems, which are typically trained on monolingual data and thus struggle to handle the intricacies of code-switching across different linguistic levels, such as syntax, semantics, and phonetics [16, 17].
The Dravidian languages, with their rich cultural and linguistic history, present a distinct case for code-mixed text analysis [18, 19]. Tamil, with its official status in India, Sri Lanka, and Singapore, and a significant diaspora community, has a script that evolved from ancient scripts like Tamili and the Chola-Pallava script [20, 21, 22]. Similarly, Malayalam, predominantly spoken in Kerala, India, employs an alpha-syllabic script that blends alphabetic and syllable-based writing systems [23, 24, 25]. However, on social media, these languages are often typed using the Roman script, contributing to a growing corpus of code-mixed data.
To address these challenges, this paper introduces a prompt-based method [12, 26, 27] using GPT-3.5 Turbo [28, 29] for sarcasm and sentiment detection in code-mixed texts, specifically focusing on Tamil-English and Malayalam-English language pairs. This shared task [30, 31, 32, 33, 34, 35, 36, 37, 38, 39] aims to advance research in sentiment analysis by providing a dataset that reflects real-world scenarios where code-mixing is prevalent and by encouraging the development of systems that can accurately classify sarcasm and sentiment polarity in these under-resourced languages.
@@
In this work, we explore the capabilities of state-of-the-art large language models, specifically GPT-3.5 Turbo [28, 39], by leveraging prompt-based techniques to classify social media comments into sarcastic or non-sarcastic categories. GPT-3.5 Turbo [28], a variant of OpenAI’s GPT-3.5 model, is known for its advanced natural language understanding and generation capabilities, making it a promising candidate for handling the nuanced task of sarcasm detection, particularly in code-mixed texts.
Our team, TextTitans, participated in a competition aimed at advancing sarcasm detection in Tamil and Malayalam. The results of our systems’ performance provide valuable insights into the current state of NLP for these languages. For Tamil, our system achieved a macro-F1 score of 0.61, placing us 9th among competing teams. The macro-F1 score, which balances precision and recall across multiple classes, reflects our system’s ability to effectively handle the nuances of Tamil sarcasm. However, despite the relatively strong performance indicated by the F1 score, the 9th rank suggests that the competition was particularly fierce, with other systems either matching or surpassing our performance.
@@ 2 Related Work
Sarcasm detection has been a topic of significant interest in the field of natural language processing (NLP) due to its complex and context-dependent nature [40]. Traditional approaches to sarcasm detection primarily relied on feature-based methods, where linguistic cues such as punctuation, lexical patterns, and syntactic structures were used to identify sarcasm [40]. For instance, works by [41, 42] and [43] utilized syntactic patterns and hashtag-based supervision to improve sarcasm detection in English texts. These approaches, however, often struggled with generalization, particularly in scenarios involving nuanced and culturally specific forms of sarcasm.
[44] introduced an approach using deep neural networks that leveraged context to improve sarcasm detection. Similarly, [45, 46] explored attention-based models that better captured the dependencies within a sentence that could indicate sarcasm. These deep learning models significantly advanced the state-of-the-art in sarcasm detection, particularly in monolingual contexts.
One of the primary challenges in dealing with code-mixed text is language identification [47, 48, 49], which involves determining the language of each word or phrase within a mixed-language sentence. Traditional methods for this task often relied on n-gram models, dictionary-based approaches, and supervised learning techniques. However, these methods sometimes struggle with the complex linguistic patterns present in code-mixed data. These models are particularly effective at capturing the contextual information necessary to accurately identify languages in code-mixed text.
Another critical area of research is part-of-speech tagging [50, 51] for code-mixed sentences. This task becomes more complicated due to the interaction between different languages’ grammatical structures. Researchers have employed techniques like Conditional Random Fields (CRFs) and sequence-to-sequence models, adapting them to handle the intricacies of mixed-language data. The use of multilingual embeddings, where words from different languages are mapped into a shared vector space, has also been shown to improve the accuracy of part-of-speech tagging in code-mixed scenarios.
@@
Additionally, speech recognition [54, 55] in code-mixed environments requires models that can accurately transcribe spoken language that switches between different languages. This task is particularly challenging because traditional speech recognition systems are typically designed for single-language input. Recent advances involve the use of end-to-end models, such as attention-based encoder-decoder architectures, which have shown promise in recognizing and transcribing code-mixed speech.
@@ 5 Methodology
  • Reduced Need for Task-Specific Data: Building a model from scratch or fine-tuning a model for sarcasm detection in code-mixed texts would typically require a substantial amount of labeled data, which is often difficult to obtain for under-resourced languages [61]. Prompting, however, minimizes the need for such extensive labeled datasets, as it allows the model to utilize its general language understanding capabilities with minimal task-specific adjustments [26, 62]. This makes it a more feasible and efficient solution for resource-constrained settings.
@@ 5.1 Prompt Engineering-Based Approach
We used the GPT-3.5 Turbo model via prompting to solve the classification task. We used GPT-3.5 Turbo in zero-Shot mode via prompting. After the prompt is provided to the LLM, the following steps happen internal to the LLM while generating the output. The following outlines the steps that occur internally within the LLM, summarizing the prompting approach using GPT-3.5 Turbo:
@@
Figure 1: An overview of GPT-3.5 Turbo sarcasm detection in Tamil.
@@
Figure 2: An overview of GPT-3.5 Turbo sarcasm detection in Malayalam.
@@
Similarly, for Malayalam, the prompt used was: “Please Check whether the comment-<Text> is Sarcastic or Non-sarcastic. Only state Sarcastic or Non-sarcastic”. This prompt, while similar in structure to the Tamil prompt, was designed specifically for Malayalam to maintain consistency in the classification task across both languages while accounting for the unique linguistic and cultural nuances of Malayalam.
@@ 6 Results
Our team named TextTitans, for the Tamil language the macro-F1 score is 0.61 with a rank of 9th. For Malayalam language the macro-F1 score is 0.50 with a rank of 13th.

### v2

@@ 1 Introduction
Sarcasm, a form of expression where the intended meaning sharply diverges from the literal meaning, poses a significant challenge for sentiment analysis systems [1]. This complexity arises because sarcasm often involves subtle linguistic cues and context-dependent interpretation, making it difficult for automated systems to accurately detect and classify [2, 3].
Sarcasm is a sharp, often humorous form of verbal irony where someone says the opposite of what they truly mean, typically to convey disdain or mockery. It relies heavily on tone and context, as the delivery can transform a seemingly straightforward statement into a biting comment[2, 3]. While it can be a playful way to highlight absurdities or critique behavior, sarcasm can also lead to misunderstandings, especially in written communication where vocal inflections are absent. Ultimately, it’s a double-edged sword that can entertain or offend, depending on the audience and the intent behind it.
@@
Code-mixing [4] in NLP refers to the blending of two or more languages within a single sentence or conversation, commonly seen in multilingual societies. It poses unique challenges for tasks like language identification, part-of-speech tagging, and sentiment analysis. Traditional NLP models often struggle with code-mixed data due to the complex interaction between languages. To address this, researchers have developed advanced neural network models, such as RNNs and CNNs, that better capture linguistic context. Code-mixed text also complicates machine translation and speech recognition, requiring specialized approaches to handle language switching effectively. As multilingualism becomes more prevalent, the importance of developing robust NLP systems capable of processing code-mixed data continues to grow.
Code-mixing is a widespread phenomenon in multilingual societies, particularly in Dravidian languages such as Tamil and Malayalam, where users frequently switch between native languages and English [4]. These code-mixed texts often appear in non-native scripts, such as Roman, due to the convenience of typing on digital platforms [4]. This creates a unique set of challenges for natural language processing (NLP) systems, which are typically trained on monolingual data and thus struggle to handle the intricacies of code-switching across different linguistic levels, such as syntax, semantics, and phonetics [5].
The Dravidian languages, with their rich cultural and linguistic history, present a distinct case for code-mixed text analysis [6]. Tamil, with its official status in India, Sri Lanka, and Singapore, and a significant diaspora community, has a script that evolved from ancient scripts like Tamili and the Chola-Pallava script [7]. Similarly, Malayalam, predominantly spoken in Kerala, India, employs an alpha-syllabic script that blends alphabetic and syllable-based writing systems [8]. However, on social media, these languages are often typed using the Roman script, contributing to a growing corpus of code-mixed data.
To address these challenges, this paper introduces a prompt-based method [9] using GPT-3.5 Turbo [10] for sarcasm and sentiment detection in code-mixed texts, specifically focusing on Tamil-English and Malayalam-English language pairs. This shared task [11, 12, 13, 14, 15, 16] aims to advance research in sentiment analysis by providing a dataset that reflects real-world scenarios where code-mixing is prevalent and by encouraging the development of systems that can accurately classify sarcasm and sentiment polarity in these under-resourced languages.
@@
In this work, we explore the capabilities of state-of-the-art large language models, specifically GPT-3.5 Turbo [10], by leveraging prompt-based techniques to classify social media comments into sarcastic or non-sarcastic categories. GPT-3.5 Turbo [10], a variant of OpenAI’s GPT-3.5 model, is known for its advanced natural language understanding and generation capabilities, making it a promising candidate for handling the nuanced task of sarcasm detection, particularly in code-mixed texts.
Our team, TextTitans, participated in a competition aimed at advancing sarcasm detection in Tamil and Malayalam. The results of our systems’ performance provide valuable insights into the current state of NLP for these languages. For Tamil, our system achieved a macro-F1 score of 0.61, placing us 9th among competing teams. The macro-F1 score, which balances precision and recall across multiple classes, reflects our system’s ability to effectively handle the nuances of Tamil sarcasm. However, despite the relatively strong performance indicated by the F1 score, we achieved 9th rank.
@@ 2 Related Work
Sarcasm detection has been a topic of significant interest in the field of natural language processing (NLP) due to its complex and context-dependent nature [17]. Traditional approaches to sarcasm detection primarily relied on feature-based methods, where linguistic cues such as punctuation, lexical patterns, and syntactic structures were used to identify sarcasm [17]. For instance, works by [18] and [19] utilized syntactic patterns and hashtag-based supervision to improve sarcasm detection in English texts. These approaches, however, often struggled with generalization, particularly in scenarios involving nuanced and culturally specific forms of sarcasm.
[20] introduced an approach using deep neural networks that leveraged context to improve sarcasm detection. Similarly, [21] explored attention-based models that better captured the dependencies within a sentence that could indicate sarcasm. These deep learning models significantly advanced the state-of-the-art in sarcasm detection, particularly in monolingual contexts.
One of the primary challenges in dealing with code-mixed text is language identification [22], which involves determining the language of each word or phrase within a mixed-language sentence. Traditional methods for this task often relied on n-gram models, dictionary-based approaches, and supervised learning techniques. However, these methods sometimes struggle with the complex linguistic patterns present in code-mixed data. These models are particularly effective at capturing the contextual information necessary to accurately identify languages in code-mixed text.
Another critical area of research is part-of-speech tagging [23] for code-mixed sentences. This task becomes more complicated due to the interaction between different languages’ grammatical structures. Researchers have employed techniques like Conditional Random Fields (CRFs) and sequence-to-sequence models, adapting them to handle the intricacies of mixed-language data. The use of multilingual embeddings, where words from different languages are mapped into a shared vector space, has also been shown to improve the accuracy of part-of-speech tagging in code-mixed scenarios.
@@
Additionally, speech recognition [26] in code-mixed environments requires models that can accurately transcribe spoken language that switches between different languages. This task is particularly challenging because traditional speech recognition systems are typically designed for single-language input. Recent advances involve the use of end-to-end models, such as attention-based encoder-decoder architectures, which have shown promise in recognizing and transcribing code-mixed speech.
@@ 5 Methodology
  • Reduced Need for Task-Specific Data: Building a model from scratch or fine-tuning a model for sarcasm detection in code-mixed texts would typically require a substantial amount of labeled data, which is often difficult to obtain for under-resourced languages [32]. Prompting, however, minimizes the need for such extensive labeled datasets, as it allows the model to utilize its general language understanding capabilities with minimal task-specific adjustments [9]. This makes it a more feasible and efficient solution for resource-constrained settings.
@@ 5.1 Prompt Engineering-Based Approach
We used the GPT-3.5 Turbo model via prompting to solve the classification task. We used GPT-3.5 Turbo in zero-Shot mode via prompting. After the prompt is provided to the LLM, the following steps happen internal to the LLM while generating the output. The following outlines the steps [10, 34, 35] that occur internally within the LLM, summarizing the prompting approach using GPT-3.5 Turbo:
@@
Figure 1: Architechture diagram for sarcasm detection in Tamil.
@@
Figure 2: Architechture diagram for sarcasm detection in Malayalam.
@@
We have used two different figures (Figure 1 and Figure 2) for the two sarcasm detection models in Tamil and Malayalam.
@@
Similarly, for Malayalam, the prompt used was: “Please Check whether the comment-<Text> is Sarcastic or Non-sarcastic. Only state Sarcastic or Non-sarcastic”. This prompt, is similar in structure to the Tamil prompt, was designed to maintain consistency in the classification task across both languages.
@@ 6 Results
Our team named TextTitans [36], for the Tamil language the macro-F1 score is 0.61 with a rank of 9th. For Malayalam language the macro-F1 score is 0.50 with a rank of 13th.
@@
Error Analysis: According to Table 1, the sarcasm detection results for the Malayalam language using GPT-3.5 reveal significant performance gaps, particularly in identifying sarcastic remarks. The model’s low precision of 0.18 for the sarcastic class indicates a high rate of false positives, suggesting that non-sarcastic comments are often misclassified as sarcastic. Additionally, a recall of 0.27 shows that many actual sarcastic instances go undetected, resulting in a low F1-score of 0.22. In contrast, the non-sarcastic class performs better, with a precision of 0.82 and an F1-score of 0.77, reflecting the model’s stronger ability to identify straightforward statements. This imbalance highlights the model’s struggle with the nuances of sarcasm, indicating the need for further refinement in training strategies or additional data to enhance performance in Malayalam.
Similarly, the sarcasm detection results for the Tamil language (See Table 2) indicate critical areas for improvement. The model’s precision for sarcastic comments is only 0.43, which signifies a considerable number of false positives, where non-sarcastic statements are misclassified as sarcastic. The recall for the sarcastic class is also low at 0.43, suggesting that the model fails to identify many genuine sarcastic instances, leading to an F1-score of 0.43. Conversely, the non-sarcastic class shows stronger performance, achieving a precision, recall, and F1-score of 0.79. This indicates effective detection of straightforward statements. The overall balanced performance, with a macro average F1-score of 0.61, highlights the inconsistency in sarcasm detection. These findings underscore the necessity for improved training methods or more diverse datasets to bolster the model’s capability to recognize sarcasm in Tamil more effectively.